The enzyme tyrosinase is an oxidase. It oxidizes phenols and catechols into quinones (aromatics with two carbonyl groups), producing brown-colored products.

In mammals, the enzyme is responsible for producing melanin in the skin and thus darkening the skin color.In plants, the enzyme is responsible for the browning that occurs when apples or bananas are opened to air.Food scientists have found that the browning reaction of tyrosinase can be prevented by adding dodecyl gallate to foods.

Using the Lineweaver-Burk plot below, calculate the Vmax and K for the enzyme in the presence and absence of the inhibitor.(The top line (squares and red) is with the inhibitor; the bottom line (diamonds and blue) is without the inhibitor). What type of inhibitor is dodecyl gallate to tyrosinase?Explain.

10г y-1.58x + 4.27 8 6 S- 2 y 1.52x1.51 1S id 1.0 2.0 3.0 -3.0 2.0 0.0 1/IS] (mM-1) ir
Answer #1

1) V max & K m in presence of inhibitor

The equation of line is y=1.58 x + 4.27 Comparing to y=mx+c we get slope= 1.58 and intercept= 4.27.

1/ V max= 4.27

Therefore, V max=1/4.27= 0.2342 OD-1min

Slope = Km / V max

Km= slope * Vmax= 1.58 * 0.2342= 0.37 mM-1

2) V max & K m in absence of inhibitor

The equation of line is y=1.52 x + 1.51 Comparing to y=mx+c we get slope= 1.52 and intercept= 1.51

1/ V max= 1.51

Therefore, V max=1/1.51= 0.6623 OD-1min

Slope = Km / V max

Km= slope * Vmax= 1.52* 0.6623= 1.006  mM-1